At the center of a 50-m-diameter circular ice rink, a 79 kg skater traveling north at 2.9 m/s collides with and holds onto a 65 kg skater who had been heading west at 3.3 m/s .

How long will it take them to glide to the edge of the rink?

Where will they reach it? Give your answer as an angle north of west.
Answer #1

Here ,

for the final speed

north is the positive y direction

west is the positive x direction

(79 + 65) * v = 79 * 2.9 j + 65 * 3.3 * i

solving for v

v = 1.59 j + 1.49 i

v = 2.18 m/s at 47 degree

for the time taken = 25/2.18 = 11.5 s

the time taken for them to reach the edge is 11.5 s

they will reach at 47 degree north of west